OpenStudy (anonymous):

find all values of Θ for which sinΘ = -sqrt (1-cos^2Θ) on (0,2pi)

OpenStudy (anonymous):

the fundamental law that is [sin²x + cos²x = 1] will help you that is the same as sin²x=1-cos²x

OpenStudy (anonymous):

\[\sin ^{2}\theta + \cos ^{2}\theta = 1\] Very important formula